OpenStudy (anonymous):

Find the equation of a parabola with vertex on the line y=x, axis parallel to 0x, and passing through (6,-2),(3,4).

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

never seen an equation like this

geerky42 (geerky42):

wow tough question.

geerky42 (geerky42):

give me min

OpenStudy (anonymous):

this is probably my assignment

OpenStudy (anonymous):

and i have exam tommorow lol

geerky42 (geerky42):

I don't think such parabola exists.

geerky42 (geerky42):

https://www.desmos.com/calculator/ivqqpnqxih

geerky42 (geerky42):

Do you have answer?

OpenStudy (anonymous):

i have its \[y^2-4x-4y+12=0: 5y^2-36x-28y+140=0\]

geerky42 (geerky42):

huh, problem said axis of symmetry is parallel to x=0?

OpenStudy (anonymous):

can you help me to factor this, \[55x^2-72x+84\]

geerky42 (geerky42):

so parabola must be vertical? I am confused

OpenStudy (anonymous):

do you know site that can factor that out?

OpenStudy (anonymous):

or its not factorable?

geerky42 (geerky42):

It's not factorable

OpenStudy (anonymous):

Let A( h, h ) be the vertex. Since the axis of symmetry is parallel to OX, the equation of this parabola is ( y - h )² = 4a ( x - h ) .................... (1) ......................................... Since it passes through the points ( 6, -2 ) and ( 3, 4 ), (-2-h)² = 4a(6-h) ........... (2) (4-h)² = 4a(3-h) ............. (3) Dividing EQ(1) by EQ(2), (2+h)² / (4-h)² = (6-h) / (3-h). Solving this will give you h. Put this value of h in (2) to get a. Then, put these values of a and h in (1) to get the final answer.

geerky42 (geerky42):

Because it has no real root

OpenStudy (anonymous):

its from google

OpenStudy (anonymous):

Dividing EQ(1) by EQ(2), (2+h)² / (4-h)² = (6-h) / (3-h). how to work with this?

geerky42 (geerky42):

Exactly what does OX mean?

OpenStudy (anonymous):

IDK, D:

geerky42 (geerky42):

Not really sure. Your problems is kind of abusive.

geerky42 (geerky42):

But I would expand everything, then multiply both sides by denominator, then combine like terms.

geerky42 (geerky42):

ah I get it.

OpenStudy (anonymous):

how?

geerky42 (geerky42):

gee I wrote a lot of LaTeX and Openstudy crashes.. all time wasted

geerky42 (geerky42):

I'm just write on paper then post it, is that ok to you?

OpenStudy (anonymous):

yup,i will appreciate it more if you do :D

geerky42 (geerky42):

ok tahes me time, is that ok?

geerky42 (geerky42):

it would takes me time*

OpenStudy (anonymous):

if you write it, can you try this problem too... find the point of intersection of the given curves and draw the figure \[x^2= 4y, x^2=y+3\]

geerky42 (geerky42):

Ok, there is good chance I won't be here when you wake up, so I seriously pray these papers make a difference. For \(\dfrac{(2+h)^2}{(4-h)^2}=\dfrac{6-h}{3-h}\):

geerky42 (geerky42):

And for point of intersection problem:

geerky42 (geerky42):

grr forgot to rotate photos. so here are them again:

geerky42 (geerky42):

For figures from same problem: https://www.desmos.com/calculator/qbsw1ybtnw